# Boundary conditions matter: On the spectrum of infinite quantum graphs

### v2

@@ 4.2.2. Inradius
$\operatorname{\operatorname{Inr}}(\mathcal{G},\mathfrak{V}):=\sup\{\operatorname{dist}_{\mathcal{G}}(x,\mathfrak{D}):x\in\mathcal{G}\},$
@@
We say that $\mathsf{v}_{\mathsf{c}}$ is a centre vertex of $\mathcal{G}$ if $\operatorname{dist}(\mathsf{v}_{\mathsf{c}},\gamma_{1})=\operatorname{dist}(\mathsf{v}_{\mathsf{c}},\gamma_{2})$ holds for any $\gamma_{1},\gamma_{2}\in\mathfrak{V}$. In that case we call $\mathcal{G}$ a centred tree.
In particular, in the trees we consider, there is at most one vertex of degree one (namely $\mathsf{v}_{\mathsf{c}}$, if it is of degree one) where standard vertex conditions are imposed, whereas on all all other vertices, we impose Dirichlet conditions. Note that, if $\mathsf{v}_{\mathsf{c}}$ is a centre vertex of $\mathcal{G}$, then
@@
Let $\mathsf{E}_{1}$ denote the edge set of $\mathcal{G}_{1}$ and let $\psi_{1}$ denote the restriction of $\psi$ to $\mathcal{G}_{1}$. For a given vertex $\mathsf{v}\in\mathcal{G}_{1}\cap\mathcal{G}_{2}$ let $\mathsf{E}_{2,\mathsf{v}}$ denote the set of edges in $\mathcal{G}_{2}$ that are incident to $\mathsf{v}$, and suppose that for each such edge $\mathsf{e}\simeq[0,\ell_{\mathsf{e}}]$ the vertex $\mathsf{v}$ is identified with 00. Then, our assumption states that $\psi^{\prime}_{\mathsf{e}}(0)\leq 0$ for all $\mathsf{e}\in\mathsf{E}_{2,\mathsf{v}}$. A direct calculation using integration by parts yields
@@
Step 2: Suppose now that $\mathcal{G}$ is a compact tree and $\mathsf{v}_{\mathsf{c}}$ has degree $d>1$. Again let $\psi$ denote a nonnegative eigenfunction corresponding to $\lambda_{1}(\mathcal{G},\mathfrak{V})$. Since $\psi$ satisfies Kirchhoff conditions in $\mathsf{v}_{\mathsf{c}}$, there exists an edge $\mathsf{e}$ incident to $\mathsf{v}_{\mathsf{c}}$ such that $\psi$ has nonpositive derivative on $\mathsf{e}$ at $\mathsf{v}_{\mathsf{c}}$. Now, cutting through the centre vertex $d-1$ times yields $d$ disjoint trees; for each of these, $\mathsf{v}_{\mathsf{c}}$ continues to be the centre vertex (and in particular each has the same inradius). From these trees, let $\mathcal{G}_{2}$ denote the tree containing the edge $\mathsf{e}$ and let $\mathcal{G}_{1}$ denote its complement in $\mathcal{G}$, the (restored) union of the other $d-1$ trees. In $\mathcal{G}_{1}$ the centre vertex $\mathsf{v}_{\mathsf{c}}$ has degree $d-1$. Setting $\mathfrak{V}_{1}:=\mathfrak{V}\cap\mathcal{G}_{1}$ and applying Lemma 4.19 we obtain $\lambda_{1}(\mathcal{G},\mathfrak{V})\geq\lambda_{1}(\mathcal{G}_{1},\mathfrak{V}_{1})$. An induction argument together with Step 1 now yields (4.19) for all compact trees.
@@
Indeed, call $\mathcal{G}_{T}^{+}$ the graph obtained by adding another edge of length $\ell_{2}$ to $\mathsf{v}$ and call $\mathfrak{V}^{+}$ the set of its degree one vertices. The graph $\mathcal{G}_{T}^{+}$ can be understood as two Dirichlet intervals $[0,\ell_{1}+\ell_{2}]$, glued at one point. By symmetry, the restriction of any ground state $\psi$ (i.e. eigenfunction for $\lambda_{1}(\mathcal{G}_{T},\mathfrak{V})$) must coincide with the corresponding Dirichlet ground states on each interval. On the one hand, this implies
@@ 5. Upper bounds on the eigenvalues
The proof follows directly from combining [36, Theorem 7.1] (which gives (5.1) on any compact graph) applied to any compact exhaustion of $\mathcal{G}$, together with Proposition 2.11 and Lemma 3.8(3). ∎
@@
In the case of a compact tree graph (i.e., for which $\beta=0$) we recover a theorem of Rohleder [56, Theorem 3.4]; for general compact graphs it improves the upper bound on $\mu_{2}$ in [36, Remark 6.3], $\mu_{2}(\mathcal{G})\leq\frac{4\pi^{2}|\mathcal{E}|^{2}}{D^{2}}$, and provides a corresponding bound on the higher eigenvalues $\mu_{k}$, $k\geq 3$, for the first time.
@@
Since $\mathcal{G}_{n}$ has Betti number $\beta$, it is possible to cut it $\beta$ times to produce a tree $\mathcal{T}$, which obviously satisfies $\operatorname{Diam}(\mathcal{T})\geq\operatorname{Diam}(\mathcal{G}_{n})$. By Lemma 4.9, we have $\mu_{k}(\mathcal{G}_{n})\leq\mu_{k+\beta}(\mathcal{T})$. Now [56, Theorem 3.4] implies that
@@
If finite length of $\mathcal{G}$ is assumed, then for girth to be a nontrivial quantity there can only be finitely many cycles (and ends) in $\mathcal{G}$, since otherwise the infimum will necessarily be zero. In this case, we can immediately reproduce the bound [11, Proposition 1.8].
@@
The proof requires adapting to infinite graphs the construction of a suitable test function as done for finite graphs in the proof of [11, Proposition 1.8]: we omit the easy details.
Other bounds, such as [11, Theorem 1.2], stating that in the presence of at least one Dirichlet vertex,
@@ Appendix A Relation of the Freudenthal compactification to the metric completion
[…]
@@
for all $n\geq n_{0}$. Since $\partial U_{n_{0}}$ and $\partial U_{n_{0}}^{\prime}$ are compact and disjoint, we have $\operatorname{dist}_{\mathcal{G}}(\partial{U_{n_{0}}},\partial{U_{n_{0}}^{\prime}})>0$. Thus, (A.3) is a contradiction to $(x_{n})_{n\in\mathbb{N}}$ and $(x_{n}^{\prime})_{n\in\mathbb{N}}$ having the same limit in $\overline{\mathcal{G}}$. In conclusion, $U_{n}\cap U_{n}^{\prime}$ must be non-empty for all $n\in\mathbb{N}$. Next, we prove $\gamma=\gamma^{\prime}$. Let $n\in\mathbb{N}$ be fixed. Since $\partial U_{n}$ is compact and the open sets $(\mathcal{G}\setminus\overline{U_{j}^{\prime}})_{j\in\mathbb{N}}$ exhaust $\mathcal{G}$, there exists some $j\geq n$ with $\partial U_{n}\subset\mathcal{G}\setminus U_{j}^{\prime}$. This implies that
@@
holds. Note that $U_{j}^{\prime}\cap U_{n}$ is a superset of the non-empty set $U_{j}^{\prime}\cap U_{j}$ and is thus nonempty as well. Therefore, since both sets on the right-hand side of (A.4) are open and $U_{j}^{\prime}$ is connected, we obtain $U_{j}^{\prime}\cap(\mathcal{G}\setminus\overline{U_{n}})=\emptyset$ and, thus, $U_{j}^{\prime}\subset U_{n}$. Analogously, one shows the existence of $k\in\mathbb{N}$ with $U_{k}\subset U_{j}^{\prime}$. Therefore, $(U_{n})_{n\in\mathbb{N}}$ and $(U^{\prime}_{n})_{n\in\mathbb{N}}$ are equivalent with respect to the relation from Definition 2.4, which in turn yields $\gamma=\gamma^{\prime}$. ∎
@@
We first prove the implication (i) $\implies$ (ii). By definition of $\beta_{\mathcal{G}}$, $\beta_{\mathcal{G}}(x)=\gamma$ yields that there is a representative $(V_{k})_{k\in\mathbb{N}}$ of $\gamma$ with $x_{k}\in V_{k}$ for all $k\in\mathbb{N}$. By equivalence, for $n\in\mathbb{N}$, there exists $k_{n}\in\mathbb{N}$ with $V_{k_{n}}\subset U_{n}$. For $k\geq k_{n}$ we obtain $x_{k}\in V_{k}\subset V_{k_{n}}\subset U_{n}$. This yields (ii). For the implication (ii) $\implies$ (iii), suppose (iii) does not hold. Then, for some $n\in\mathbb{N}$, we find a sequence $(x_{k})_{k\in\mathbb{N}}$ with $x_{k}\ni U_{n}$ and $\operatorname{dist}_{\mathcal{G}}(x,x_{k})<\frac{1}{k}$ for all $k\in\mathbb{N}$. Therefore, $(x_{k})_{k\in\mathbb{N}}$ is a Cauchy sequence converging to $x$ that does not satisfy (ii). By contraposition (ii) implies (iii). It remains to prove that (iii) implies (i). By (iii), for $n\in\mathbb{N}$, there exists some $\varepsilon_{n}>0$ with
@@
[…]
@@
We consider the infinite ladder graph $\mathcal{G}$ as depicted in Figure A.1. The ladder graph has exactly one topological end $\gamma$. However, hereinafter, we consider four different choices for the lengths of the graph edges $\mathsf{e}_{n},\mathsf{f}_{n}^{1}$ and $\mathsf{f}_{n}^{2}$ for $n\in\mathbb{N}$ (see Figure A.1 for the notation), each of which leads to a different structure of the metric boundary of $\mathcal{G}$.
@@
[…]
@@
is an neighbourhood of the unique point $x\in\overline{\mathcal{G}}\setminus\mathcal{G}$, whereas its image
@@
$\mathsf{f}_{1}^{1}$$\mathsf{f}_{2}^{1}$$\mathsf{f}_{3}^{1}$$\mathsf{f}_{1}^{2}$$\mathsf{f}_{2}^{2}$$\mathsf{f}_{3}^{2}$$\mathsf{e}_{1}$$\mathsf{e}_{2}$$\mathsf{e}_{3}$$\mathsf{e}_{4}$
Figure A.1. The infinite ladder graph.
@@
the limit $x$ of $(x_{n})_{n\in\mathbb{N}}$ must be in $\overline{\mathcal{G}}\setminus\mathcal{G}$. Moreover, by definition of $\beta_{\mathcal{G}}$ we have $\beta_{\mathcal{G}}(x)=\gamma$. It remains to show the uniqueness of $x$. Let $x,y\in\overline{\mathcal{G}}\setminus\mathcal{G}$ with $\beta_{\mathcal{G}}(x)=\beta_{\mathcal{G}}(y)=\gamma$. Let $(U_{n})_{n\in\mathbb{N}}$ be a representative of $\gamma$ and let $(x_{n})_{n\in\mathbb{N}}$ and $(y_{n})_{n\in\mathbb{N}}$ be two Cauchy sequences in $\mathcal{G}$ converging to $x$ and $y$ respectively. Let $n\in\mathbb{N}$. By Lemma A.5(ii), there exists some $k_{n}\in\mathbb{N}$ with $x_{k_{n}},y_{k_{n}}\in U_{n}$. We obtain
@@
An graph end of infinite volume $\gamma$ of $\mathcal{G}$ with $\beta_{\mathcal{G}}^{-1}(\gamma)\neq\emptyset$ is called a degenerate graph end.
@@
We claim that there exists some $n\in\mathbb{N}$ with $U_{n}\subset U_{\varepsilon}(x)$. Assume for a contradiction that for each $n\in\mathbb{N}$, there exists some $y_{n}\in U_{n}$ with $\operatorname{dist}_{\mathcal{G}}(x,y_{n})\geq\varepsilon$. As in the proof of Proposition A.8, it follows that $(y_{n})_{n\in\mathbb{N}}$ converges to some $y\in\overline{\mathcal{G}}\setminus\mathcal{G}$ with $\beta_{\mathcal{G}}(y)=\gamma$. Moreover, by construction of $(y_{n})_{n\in\mathbb{N}}$ we have $\operatorname{dist}_{\mathcal{G}}(x,y)\geq\varepsilon$ and, thus, $x\neq y$. However, by assumption, $\gamma$ must be a finite graph end. Therefor, $x\neq y$ contradicts Proposition A.8. We conclude that there exists some $n\in\mathbb{N}$ with $U_{n}\subset U_{\varepsilon}(x)\subset\beta_{\mathcal{G}}(W)$. This completes the proof. ∎

### v3

@@ 4.2.2. Inradius
$\operatorname{\operatorname{Inr}}(\mathcal{G},\mathfrak{V}):=\sup\{\operatorname{dist}_{\mathcal{G}}(x,\mathfrak{V}):x\in\mathcal{G}\},$
@@
We say that $\mathsf{v}_{\mathsf{c}}$ is a centre vertex of $\mathcal{G}$ if $\operatorname{dist}(\mathsf{v}_{\mathsf{c}},\gamma_{1})=\operatorname{dist}(\mathsf{v}_{\mathsf{c}},\gamma_{2})$ holds for any $\gamma_{1},\gamma_{2}\in\mathfrak{V}$. Whenever (at least) a centre vertex exists, we call $\mathcal{G}$ a centred tree.
We refer to Figure 3.1 for an example of a metric tree with two centre vertices. In particular, in the trees we consider, there is at most one vertex of degree one (namely $\mathsf{v}_{\mathsf{c}}$, if it is of degree one) where standard vertex conditions are imposed, whereas on all all other vertices, we impose Dirichlet conditions. Note that, if $\mathsf{v}_{\mathsf{c}}$ is a centre vertex of $\mathcal{G}$, then
@@
Figure 4.2 visually describes a possible case of application of Lemma 4.19.
@@
Let $\mathsf{E}_{1}$ denote the edge set of $\mathcal{G}_{1}$ and let $\psi_{1}$ denote the restriction of $\psi$ to $\mathcal{G}_{1}$. For a given vertex $\mathsf{v}\in\mathcal{G}_{1}\cap\mathcal{G}_{2}$ let $\mathsf{E}_{2,\mathsf{v}}$ denote the set of edges in $\mathcal{G}_{2}$ that are incident to $\mathsf{v}$, and suppose that for each such edge $\mathsf{e}\simeq[0,\ell_{\mathsf{e}}]$ the vertex $\mathsf{v}$ is identified with 0. Then, our assumption states that $\psi^{\prime}_{\mathsf{e}}(0)\leq 0$ for all $\mathsf{e}\in\mathsf{E}_{2,\mathsf{v}}$. A direct calculation using integration by parts yields
@@
Step 2: Suppose now that $\mathcal{G}$ is a compact tree and $\mathsf{v}_{\mathsf{c}}$ has degree $d>1$. Again let $\psi$ denote a nonnegative eigenfunction corresponding to $\lambda_{1}(\mathcal{G},\mathfrak{V})$. Since $\psi$ satisfies Kirchhoff conditions in $\mathsf{v}_{\mathsf{c}}$, there exists an edge $\mathsf{e}$ incident to $\mathsf{v}_{\mathsf{c}}$ such that $\psi$ has nonpositive derivative on $\mathsf{e}$ at $\mathsf{v}_{\mathsf{c}}$. Now, cutting through the centre vertex d1d-1 times yields $d$ disjoint trees; for each of these, $\mathsf{v}_{\mathsf{c}}$ continues to be the centre vertex (and in particular each has the same inradius). From these trees, let $\mathcal{G}_{2}$ denote the tree containing the edge $\mathsf{e}$ and let $\mathcal{G}_{1}$ denote its complement in $\mathcal{G}$, the (restored) union of the other d1d-1 trees. In $\mathcal{G}_{1}$ the centre vertex $\mathsf{v}_{\mathsf{c}}$ has degree d1d-1. Setting $\mathfrak{V}_{1}:=\mathfrak{V}\cap\mathcal{G}_{1}$ and applying Lemma 4.19 we obtain $\lambda_{1}(\mathcal{G},\mathfrak{V})\geq\lambda_{1}(\mathcal{G}_{1},\mathfrak{V}_{1})$. An induction argument together with Step 1 now yields (4.19) for all compact trees.
@@
Indeed, call $\mathcal{G}_{T}^{+}$ the graph obtained by adding another edge of length $\ell_{2}$ to $\mathsf{v}$ and call $\mathfrak{V}^{+}$ the set of its degree one vertices. The graph $\mathcal{G}_{T}^{+}$ can be understood as two Dirichlet intervals $0,\ell_{1}+\ell_{2}]$, glued at one point. By symmetry, the restriction of any ground state $\psi$ (i.e. eigenfunction for $\lambda_{1}(\mathcal{G}_{T},\mathfrak{V})$) must coincide with the corresponding Dirichlet ground states on each interval. On the one hand, this implies
@@ 5. Upper bounds on the eigenvalues
The proof follows directly from combining [41, Theorem 7.1] (which gives (5.1) on any compact graph) applied to any compact exhaustion of $\mathcal{G}$, together with Proposition 2.11 and Lemma 3.9(3). ∎
@@
In the case of a compact tree graph (i.e., for which $\beta=0$) we recover a theorem of Rohleder [60, Theorem 3.4]; for general compact graphs it improves the upper bound on $\mu_{2}$ in [41, Remark 6.3], $\mu_{2}(\mathcal{G})\leq\frac{4\pi^{2}|\mathcal{E}|^{2}}{\operatorname{Diam}(\mathcal{G})^{2}}$, and provides a corresponding bound on the higher eigenvalues $\mu_{k}$, $k\geq 3$, for the first time.
@@
Since $\mathcal{G}_{n}$ has Betti number $\beta$, it is possible to cut it $\beta$ times to produce a tree $\mathcal{T}$, which obviously satisfies $\operatorname{Diam}(\mathcal{T})\geq\operatorname{Diam}(\mathcal{G}_{n})$. By Lemma 4.9, we have $\mu_{k}(\mathcal{G}_{n})\leq\mu_{k+\beta}(\mathcal{T})$. Now [60, Theorem 3.4] implies that
@@
If finite length of $\mathcal{G}$ is assumed, then for girth to be a nontrivial quantity there can only be finitely many cycles (and ends) in $\mathcal{G}$, since otherwise the infimum will necessarily be zero. In this case, we can immediately reproduce the bound [16, Proposition 1.8].
@@
The proof requires adapting to infinite graphs the construction of a suitable test function as done for finite graphs in the proof of [16, Proposition 1.8]: we omit the easy details.
Other bounds, such as [16, Theorem 1.2], stating that in the presence of at least one Dirichlet vertex,
@@ Appendix A Relation of the Freudenthal compactification to the metric completion
Throughout this section let $\mathcal{G}$ be a locally finite metric graph. Our aim is to discuss the relation between the completion $\overline{\mathcal{G}}$ of $\mathcal{G}$ as a metric space and its Freudenthal compactification $\mathcal{G}\cup\mathfrak{C}(\mathcal{G})$, where $\mathfrak{C}(\mathcal{G})$ denotes the set of topological ends of $\mathcal{G}$ (see Definition 2.4). The following proposition shows that each point in the metric boundary $\overline{\mathcal{G}}\setminus\mathcal{G}$ can be associated with a unique end in $\mathfrak{C}(\mathcal{G})$.
@@
for all $n\geq n_{0}$. Since $\partial U_{n_{0}}$ and $\partial U_{n_{0}}^{\prime}$ are compact and disjoint, we have $\operatorname{dist}_{\mathcal{G}}(\partial{U_{n_{0}}},\partial{U_{n_{0}}^{\prime}})>0$. Thus, (A.3) is a contradiction to $(x_{n})_{n\in\mathbb{N}}$ and $(x_{n}^{\prime})_{n\in\mathbb{N}}$ having the same limit in $\overline{\mathcal{G}}$. In conclusion, $U_{n}\cap U_{n}^{\prime}$ must be non-empty for all $n\in\mathbb{N}$.
Next, we prove $\gamma=\gamma^{\prime}$. Let $n\in\mathbb{N}$ be fixed. Since $\partial U_{n}$ is compact and the open sets $(\mathcal{G}\setminus\overline{U_{j}^{\prime}})_{j\in\mathbb{N}}$ exhaust $\mathcal{G}$, there exists some $j\geq n$ with $\partial U_{n}\subset\mathcal{G}\setminus U_{j}^{\prime}$. This implies that
@@
holds. Note that $U_{j}^{\prime}\cap U_{n}$ is a superset of the non-empty set $U_{j}^{\prime}\cap U_{j}$ and is thus non-empty as well. Therefore, since both sets on the right-hand side of (A.4) are open and $U_{j}^{\prime}$ is connected, we obtain $U_{j}^{\prime}\cap(\mathcal{G}\setminus\overline{U_{n}})=\emptyset$ and, thus, $U_{j}^{\prime}\subset U_{n}$. Analogously, one shows the existence of $k\in\mathbb{N}$ with $U_{k}\subset U_{j}^{\prime}$. Therefore, $(U_{n})_{n\in\mathbb{N}}$ and $(U^{\prime}_{n})_{n\in\mathbb{N}}$ are equivalent with respect to the relation from Definition 2.4, which in turn yields $\gamma=\gamma^{\prime}$. ∎
@@
We first prove the implication (i) $\implies$ (ii). By definition of $\beta_{\mathcal{G}}$, $\beta_{\mathcal{G}}(x)=\gamma$ yields that there is a representative $(V_{k})_{k\in\mathbb{N}}$ of $\gamma$ with $x_{k}\in V_{k}$ for all $k\in\mathbb{N}$. By equivalence, for $n\in\mathbb{N}$, there exists $k_{n}\in\mathbb{N}$ with $V_{k_{n}}\subset U_{n}$. For $k\geq k_{n}$ we obtain $x_{k}\in V_{k}\subset V_{k_{n}}\subset U_{n}$. This yields (ii).
For the implication (ii) $\implies$ (iii), suppose (iii) does not hold. Then, for some $n\in\mathbb{N}$, we find a sequence $(x_{k})_{k\in\mathbb{N}}$ with $x_{k}\ni U_{n}$ and $\operatorname{dist}_{\mathcal{G}}(x,x_{k})<\frac{1}{k}$ for all $k\in\mathbb{N}$. Therefore, $(x_{k})_{k\in\mathbb{N}}$ is a Cauchy sequence converging to $x$ that does not satisfy (ii). By contraposition (ii) implies (iii).
It remains to prove that (iii) implies (i). By (iii), for $n\in\mathbb{N}$, there exists some $\varepsilon_{n}>0$ with
@@
We note that Lemma A.5(iii) implies that $\beta_{\mathcal{G}}$ is in fact a continuous map when $\overline{\mathcal{G}}$ is equipped with the topology of the completion of $\mathcal{G}$ as a metric space and $\mathcal{G}\cup\mathfrak{C}(\mathcal{G})$ is equipped with the topology of the Freudenthal compactification of $\mathcal{G}$.
@@
We consider the infinite ladder graph $\mathcal{G}$ as depicted in Figure 2.1. The ladder graph has exactly one topological end $\gamma$. However, hereinafter, we consider four different choices for the lengths of the graph edges $\mathsf{e}_{n},\mathsf{f}_{n}^{1}$ and $\mathsf{f}_{n}^{2}$ for $n\in\mathbb{N}$ (see Figure 2.1 for the notation), each of which leads to a different structure of the metric boundary of $\mathcal{G}$.
@@
We finish this example by comparing the cases (2) and (4), where $\beta_{\mathcal{G}}$ is bijective. We remark first that, in (2), it can be shown $\beta_{\mathcal{G}}$ is in fact a homeomorphism and, thus, the topologies of $\overline{\mathcal{G}}$ and $\mathcal{G}\cup\mathfrak{C}(\mathcal{G})$ coincide. However, in (4), $\beta_{\mathcal{G}}$ is not a homeomorphism. Indeed, the set
@@
is a neighbourhood of the unique point $x\in\overline{\mathcal{G}}\setminus\mathcal{G}$, whereas its image
@@
the limit $x$ of $(x_{n})_{n\in\mathbb{N}}$ must be in $\overline{\mathcal{G}}\setminus\mathcal{G}$. Moreover, by definition of $\beta_{\mathcal{G}}$ we have $\beta_{\mathcal{G}}(x)=\gamma$.
It remains to show the uniqueness of $x$. Let $x,y\in\overline{\mathcal{G}}\setminus\mathcal{G}$ with $\beta_{\mathcal{G}}(x)=\beta_{\mathcal{G}}(y)=\gamma$. Let $(U_{n})_{n\in\mathbb{N}}$ be a representative of $\gamma$ and let $(x_{n})_{n\in\mathbb{N}}$ and $(y_{n})_{n\in\mathbb{N}}$ be two Cauchy sequences in $\mathcal{G}$ converging to $x$ and $y$ respectively. Let $n\in\mathbb{N}$. By Lemma A.5(ii), there exists some $k_{n}\in\mathbb{N}$ with $x_{k_{n}},y_{k_{n}}\in U_{n}$. We obtain
@@
A graph end of infinite volume $\gamma$ of $\mathcal{G}$ with $\beta_{\mathcal{G}}^{-1}(\gamma)\neq\emptyset$ is called a degenerate graph end.
@@
We claim that there exists some $n\in\mathbb{N}$ with $U_{n}\subset U_{\varepsilon}(x)$. Assume for a contradiction that for each $n\in\mathbb{N}$, there exists some $y_{n}\in U_{n}$ with $\operatorname{dist}_{\mathcal{G}}(x,y_{n})\geq\varepsilon$. As in the proof of Proposition A.8, it follows that $(y_{n})_{n\in\mathbb{N}}$ converges to some $y\in\overline{\mathcal{G}}\setminus\mathcal{G}$ with $\beta_{\mathcal{G}}(y)=\gamma$. Moreover, by construction of $(y_{n})_{n\in\mathbb{N}}$ we have $\operatorname{dist}_{\mathcal{G}}(x,y)\geq\varepsilon$ and, thus, $x\neq y$. However, by assumption, $\gamma$ must be a finite graph end. Therefore, $x\neq y$ contradicts Proposition A.8.
We conclude that there exists some $n\in\mathbb{N}$ with $U_{n}\subset U_{\varepsilon}(x)\subset\beta_{\mathcal{G}}(W)$. This completes the proof. ∎